OpenStudy (anonymous):

Find the least common denominator for these two rational expressions x/x^2-4 -9x/x^2+3x-10

OpenStudy (anonymous):

x/(x^2 - 4) = x/(x-2)(x+2) -9x/(x^2 + 3x-10) = -9x/(x+5)(x-2) LCD = (x-2)(x+2)(x+5)